OpenStudy (anonymous):

A=p(l+r)^2 gives the amount A in dollars that P dollars will grow to in 2 yers at interest rate r (where r is given as a decimal) using compouind interest. What interest rate will cause $4000 to grow to $4326.40 in 2 years

OpenStudy (wolf1728):

Total = Principal × ( 1 + Rate )^years To solve for rate we get: log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = log(4,326.40) - log(4,000.00) ÷ 2 log(1 + rate) = 3.6361266699 - 3.6020599913 ÷ 2 log(1 + rate) = 0.0340666786 ÷ 2 log(1 + rate) = 0.0170333393 1 + rate = 10^0.0170333393 1 + rate = 1.040000000 rate = .04 Formula http://1728.org/compint2.htm Compound Interest Calculator http://1728.org/compint.htm